# Understood: Real-Time Communication Support for Adults with ADHD Using Mixed Reality

### v1

@@ 3.1. Semi-Structured Interviews
To address RQ1, we conducted semi-structured interviews with two key stakeholder groups: four licensed clinicians and $12$ adults diagnosed with ADHD. By integrating clinical expertise with firsthand lived experiences, this dual-perspective methodology enabled a holistic analysis of communication barriers and intervention efficacy, yielding multifaceted insights that directly guided system requirements and design priorities.
@@ 3.1.1. Expert Interview
We conducted semi-structured interviews with four licensed clinicians specializing in ADHD (C1–C4): two psychiatrists (C1–C2) and two therapists (C3–C4)). Participants were recruited from local hospitals and had an average of 11.75 years of experience (SD = 2.86). To be eligible, clinicians were required to have managed at least 50 ADHD cases and to have directly observed clients’ natural communication patterns. Each interview lasted approximately 60 minutes, followed IRB guidelines, and participants received a compensation of \$20 USD.
The interview protocol began with an exploration of communication challenges associated with ADHD, utilizing 15 clinically validated statements on pragmatic communication skills derived from the Diagnostic Interview for ADHD in Adults (DIVA-5) (kooij2019diva), a widely used diagnostic tool based on DSM-V criteria (american2013diagnostic). Representative items included: “Has difficulty concentrating on a conversation,” “Often changes the subject of the conversation,” and “Needs many words to express a simple idea.” All the statements can be seen at Appendix A. Clinicians quantitatively assessed the prevalence of each challenge using a 3-point Likert scale (1 = uncommonly seen, 3 = frequently seen), followed by qualitative elaboration through anonymized case examples and analysis of contextual factors influencing these communication patterns. Additionally, we conducted an open-ended evaluation of existing interventions, where participants shared the strategies they employed to support individuals with ADHD in overcoming communication challenges, discussing both their advantages and limitations.
3.1.2. Patient Interview
@@ 3.1.3. Data Analysis
To holistically map communication challenges and intervention opportunities, we conducted an integrated analysis of interview data from both clinicians and patients with ADHD.
All sessions were video-recorded, transcribed, and analyzed using affinity diagramming techniques (harboe2015real; lucero2015using), facilitated via Miro [^2] https://miro.com/app/dashboard/. To ensure analytical rigor, the research team engaged in weekly consensus meetings to interpret findings and iteratively refine thematic categories. Cross-referencing clinician insights with patient narratives on the collaborative Miro board enabled triangulation of emergent patterns, particularly behavioral manifestations of the 15 DIVA-5 communication statements observed in video recordings (e.g., topic shifts, disfluency). To validate findings, participants completed a member-checking process, reviewing preliminary results to annotate alignment or discrepancies with their experiences. This iterative feedback loop ensured fidelity to participant perspectives while refining the final analytical framework.
@@ 3.1.4. Contextual Analysis Findings
Ch1: Deficits in Working Memory. In response to #6. I have difficulty filtering and/or selecting information, all four clinicians rated this issue as 3 (frequently observed), and ten patients reported experiencing similar difficulties. Many described struggling to integrate new information with existing knowledge, often forgetting earlier content mid-conversation. This challenge was widely attributed to working memory deficits. As P3 explained, “When others are speaking, it’s hard for me to process their words. I can only remember a few of them, and I often miss important parts. I sometimes need to ask for clarification before I can respond properly.”
All twelve patients acknowledged limitations in working memory, linking them to broader communication challenges. Clinicians echoed this view. As C1 explained, “Working memory is the ability to hold and process relevant information. It’s essential for understanding context and making inferences. Deficits here contribute to communication breakdowns.” Some also noted that such deficits hindered sustained attention during conversations (#3).
Ch2: Disfluency and Pauses in Conversation. In response to #2. I have difficulty speaking fluently in long sentences, three clinicians rated this issue as 3 (frequently observed), and eight patients reported similar struggles. Two key underlying causes emerged from the interviews. The first was working memory deficits, where patients described losing track of their thoughts mid-sentence. As P5 put it, “When I try to express something, it’s like… like… (stuck)… my mind just goes blank, and I have to pause to recover.”
The second cause involved cognitive overload. Patients reported that too many thoughts surfaced simultaneously, making it difficult to organize speech. As P9 shared, “I often think of both relevant and irrelevant things at once. I need time to sort them out and find the right words, so it probably looks like I’m pausing for too long.” Several participants noted that external cues helped them resume speaking. For instance, P4 explained, “When someone guesses the word I’m stuck on, it’s like my memory kicks back in. I can keep going after that, like they’ve unblocked me.”
Ch3: Attentional Anchoring Failure Reflected by Topic Shifting. In response to #4. I often change the subject of the conversation, three clinicians rated this issue as 3 (frequently observed), and nine patients reported similar experiences. As P8 recounted, “My group was discussing a project, and suddenly I started talking about what I had for breakfast. It completely derailed the conversation.” Likewise, P10 noted, “I often bring up both relevant and irrelevant topics,” linking this to #9 (talk too much) and #11 (need a lot of words to say).
@@
F1: Limitations of Existing Intervention Methods. As noted in the Introduction, current interventions for ADHD fall into two main categories: medical and non-medical. While medical treatments such as Concerta (concerta) can have been shown to alleviate communication difficulties (maras2024effect), their side effects remain a major concern. As P2 shared, “When I took Concerta, my mind felt clearer, and communication improved. But it also made me tense, nauseous, and disrupted my sleep – so I only take it when absolutely necessary.” Given that adult ADHD is a lifelong condition, the long-term sustainability of medication use is often compromised by these adverse effects. While non-medical interventions were generally perceived as beneficial, participants highlighted their limited applicability in real-world scenarios. Strategies learned in structured environments – such as CBT or SST – often proved difficult to implement during actual conversations due to executive function deficits. As P1 explained, “In real conversations, I just don’t have the mental bandwidth to apply those strategies.” SST, in particular, was critiqued for its lack of ecological validity. C3 noted, “After socializing, patients may overanalyze their mistakes, which increases hesitation and shame in future interactions. Helping them transfer learned skills from training to real-life situations remains a key challenge.”
F2: The Need for Accessible, Therapist-Free, and Real-Time Support. Interviews highlighted a pressing need for accessible, real-time support that operates independently of professional supervision. While ADHD presents persistent communication challenges, its comparatively milder clinical presentation – relative to conditions like depression or mania – often results in less attention and fewer dedicated resources. Only 6 of the 12 participants reported ever receiving structured interventions such as CBT or SST. As P7 noted, “It’s so hard for me to book a therapy appointment. I had to find some books and hope they would help.” This lack of access underscores the value of scalable, automated solutions that provide consistent support without requiring therapist involvement. Participants also emphasized the potential of real-time, context-sensitive assistance. Unlike traditional interventions confined to clinical settings, in-situ support embedded in everyday interactions could bridge the gap between learning and application. As C4 observed, “Providing assistance on-the-fly could turn real-world environments into training spaces, helping patients build resilience and reduce stigma.” Similarly, P3 remarked, “Real-time assistance may help me apply what I’ve been learned immediately.”
@@
In addition to the major themes, several minor findings emerged from the interviews. First, regarding the issue of #12. I have difficulty waiting my turn during conversations, all four clinicians rated this behavior as 3 (frequently observed), and 11 patients reported experiencing similar challenges. However, 10 of them noted that years of social experience had helped them develop self-regulation strategies to manage impulsive behaviors, such as interrupting or speaking out of turn (#12–15). As C4 noted, “This is one of the main differences between adult and child patients. We provide less training on interrupting behaviors for adults.” Second, four participants mentioned difficulty maintaining eye contact during conversations, often due to stigma-related self-consciousness. They suggested that supportive tools could help foster more natural communication by encouraging appropriate eye contact.
@@ 3.2.1. Exploratory Brainstorming
To explore a broad range of design ideas for addressing communication challenges related to ADHD, we conducted a group brainstorming session guided by three established principles (wilson2013brainstorming): (1) focus on quantity, (2) defer judgment, and (3) encourage unconventional thinking. Participants were encouraged to contribute freely without critique, while also building upon each other’s ideas to foster creative synergy. Drawing from the prior interviews, the brainstorming session began with an exploration of commonly reported communication difficulties experienced by individuals with ADHD. Discussions addressed both the limitations of existing interventions and the specific needs for support. Participants were encouraged to generate design concepts for each identified challenge while adhering to the principle of suspending judgment and allowing unrestricted ideation. All ideas proposed during the session were systematically documented. Following the session, researchers and participants collaboratively reviewed and refined the collected concepts, synthesizing a subset of proposals with the greatest potential to address the identified challenges.
@@ 3.2.2. Design Workshop
After formulating initial, high-level design concepts, we refined and expanded these ideas through a dedicated design workshop. Storyboards served as a effective rapid visualization tool (truong2006storyboarding), offering a concrete medium to convey and deepen understanding of abstract design concepts. The workshop featured 18 storyboards (examples shown in Figure 4, with a summary at Figure 10 in Appendix B), each depicting a communication challenge, a proposed solution, or a hypothetical outcome. These storyboards were iteratively developed from the preliminary design ideas, representing concepts at varying levels of granularity.
@@ 3.2.3. Contextual Analysis Findings
F3. MR Devices Support Seamless, Polite Interaction and Real-World Skill Training and Transfer. Participants including P1, P3, and R1 considered laptops and smartphones as impractical for daily communication support, as shifting gaze between the device and conversation partner could be socially inappropriate. While teleprompters were conceptually appealing, their usability was limited by spatial constraints in typical communication settings. In contrast, MR devices such as the HoloLens [^3] https://learn.microsoft.com/en-us/hololens/ were favored for their ability to overlay information directly within the user’s field of view, allowing for more natural interactions, such as maintaining eye contact – an essential aspect of effective communication. As P6 remarked, “It’s visually intuitive to present prompts via MR while minimizing conversational disruption. I can engage with my conversation partner while seamlessly receiving information through MR interfaces – without extra effort.” Clinicians also recognized MR’s potential to bridge the gap G1 in sustainable transfer of therapeutic outcomes. As C1 explained, “MR enables SST in real-world scenarios, providing opportunities for patients to learn in context and directly apply their skills in everyday interactions.”
@@ 4.1. System Overview
Building on insights from the design workshop, we developed Understood, a MR system designed to support individuals with ADHD in proactively seeking assistance, enhancing communication fluency, and improving conversational performance. By offering interactive training in realistic settings, the system enables users to gradually refine their speech patterns, serving as a form of automated substitution therapy (DG1). Interaction with Understood is structured into three main phases (Figure 5):
@@ 4.2. Interface and Interaction Design
To accommodate the complexity of MR environments – where real-world backgrounds can introduce distractions, we adopted a minimalist and user-friendly design language. The interface employs a semi-transparent black-and-white color scheme to maintain visual clarity while minimizing cognitive load. In response to potential strain from bright backgrounds, a dark mode was implemented, featuring white text on a black background to reduce visual fatigue. Rounded corners are used consistently throughout the interface to evoke a sense of softness and approachability, ensuring the design remains unobtrusive. To ensure text accessibility, we followed established design guidelines (williams2020typeface; garnham2017typography; 10.1145/3341163.3347748) and evaluated several factors, including design consistency, legibility and readability. Based on these criteria, we selected BBC Reith (bbc_reith) as the system font. Additionally, to mitigate the lack of haptic feedback in MR environments, auditory cues were integrated to provide users with immediate and perceptible confirmation of their actions (DG5).
@@
[Uncaptioned image] Self Summarization Summarize user’s speech
[Uncaptioned image] Other Summarization Summarize partner’s speech
[Uncaptioned image] Word Suggestions Suggest subsequent words
[Uncaptioned image] Off-topic Detection Detect if speech is off-topic
[Uncaptioned image] Pop-up Animation Pop panel up on gaze

### v2

@@ 3.1. Semi-Structured Interviews
To address RQ1, we conducted semi-structured interviews with two key stakeholder groups: four licensed clinicians and 121212 adults diagnosed with ADHD. By integrating clinical expertise with firsthand lived experiences, this dual-perspective methodology enabled a holistic analysis of communication barriers and intervention efficacy, yielding multifaceted insights that directly guided system requirements and design priorities.
@@ 3.1.1. Expert Interview
We conducted semi-structured interviews with four licensed clinicians specializing in ADHD (C1–C4): two psychiatrists (C1–C2) and two therapists (C3–C4). Participants were recruited from local hospitals and had an average of 11.75 years of experience (SD = 2.86). To be eligible, clinicians were required to have managed at least 50 ADHD cases and to have directly observed clients’ natural communication patterns. Each interview lasted approximately 60 minutes, followed IRB guidelines, and participants received a compensation of \$20 USD.
The interview protocol began with an exploration of communication challenges associated with ADHD, utilizing 15 clinically validated statements on pragmatic communication skills derived from the Diagnostic Interview for ADHD in Adults (DIVA-5) (Kooij et al., 2019), a widely used diagnostic tool based on DSM-V criteria (Association et al., 2013). Representative items included: “Has difficulty concentrating on a conversation,” “Often changes the subject of the conversation,” and “Needs many words to express a simple idea.” All the statements can be seen at Appendix A. Clinicians quantitatively assessed the prevalence of each challenge using a 3-point Likert scale (1 = uncommonly seen, 3 = frequently seen), followed by qualitative elaboration through anonymized case examples and analysis of contextual factors influencing these communication patterns. Additionally, we conducted an open-ended evaluation of existing interventions, where participants shared the strategies they employed to support individuals with ADHD in overcoming communication challenges, discussing both their advantages and limitations.
3.1.2. ADHD Individuals Interview
@@ 3.1.3. Data Analysis
To holistically map communication challenges and intervention opportunities, we conducted an integrated analysis of interview data from both clinicians and individuals with ADHD.
All sessions were video-recorded, transcribed, and analyzed using affinity diagramming techniques (Harboe and Huang, 2015; Lucero, 2015), facilitated via Miro [^2] https://miro.com/app/dashboard/. To ensure analytical rigor, the research team engaged in weekly consensus meetings to interpret findings and iteratively refine thematic categories. Cross-referencing clinician insights with ADHD people narratives on the collaborative Miro board enabled triangulation of emergent patterns, particularly behavioral manifestations of the 15 DIVA-5 communication statements observed in video recordings (e.g., topic shifts, disfluency). To validate findings, participants completed a member-checking process, reviewing preliminary results to annotate alignment or discrepancies with their experiences. This iterative feedback loop ensured fidelity to participant perspectives while refining the final analytical framework.
@@ 3.1.4. Contextual Analysis Findings
[…]
Illustration summarizing three communication challenges faced by individuals with ADHD. We use the same 3 illustrations as in the teaser panel A. (1) Ch1: Deficits in Working Memory – A person appears confused with a thought bubble labeled “(BLANK),” while two others look concerned. (2) Ch2: Disfluency and Pauses in Conversation – Two individuals look uneasy; one says “um…” with a question mark above their head, suggesting a struggle to maintain conversational flow. (3) Ch3: Attentional Anchoring Failure by Topic Shifting – A person talks while pointing, with icons of a flower and soccer ball above their head, while the listener looks puzzled, indicating difficulty staying focused when the topic changes.
@@
Ch1: Deficits in Working Memory. In response to #6. I have difficulty filtering and/or selecting information, all four clinicians rated this issue as 3 (frequently observed), and ten individuals with ADHD reported experiencing similar difficulties. Many described struggling to integrate new information with existing knowledge, often forgetting earlier content mid-conversation. This challenge was widely attributed to working memory deficits. As P3 explained, “When others are speaking, it’s hard for me to process their words. I can only remember a few of them, and I often miss important parts. I sometimes need to ask for clarification before I can respond properly.”
All twelve individuals with ADHD acknowledged limitations in working memory, linking them to broader communication challenges. Clinicians echoed this view. As C1 explained, “Working memory is the ability to hold and process relevant information. It’s essential for understanding context and making inferences. Deficits here contribute to communication breakdowns.” Some also noted that such deficits hindered sustained attention during conversations (#3).
Ch2: Disfluency and Pauses in Conversation. In response to #2. I have difficulty speaking fluently in long sentences, three clinicians rated this issue as 3 (frequently observed), and eight individuals with ADHD reported similar struggles. Two key underlying causes emerged from the interviews. The first was working memory deficits, where ADHD people described losing track of their thoughts mid-sentence. As P5 put it, “When I try to express something, it’s like… like… (stuck)… my mind just goes blank, and I have to pause to recover.”
The second cause involved cognitive overload. Individuals with ADHD reported that too many thoughts surfaced simultaneously, making it difficult to organize speech. As P9 shared, “I often think of both relevant and irrelevant things at once. I need time to sort them out and find the right words, so it probably looks like I’m pausing for too long.” Several participants noted that external cues helped them resume speaking. For instance, P4 explained, “When someone guesses the word I’m stuck on, it’s like my memory kicks back in. I can keep going after that, like they’ve unblocked me.”
Ch3: Attentional Anchoring Failure Reflected by Topic Shifting. In response to #4. I often change the subject of the conversation, three clinicians rated this issue as 3 (frequently observed), and nine individuals with ADHD reported similar experiences. As P8 recounted, “My group was discussing a project, and suddenly I started talking about what I had for breakfast. It completely derailed the conversation.” Likewise, P10 noted, “I often bring up both relevant and irrelevant topics,” linking this to #9 (talk too much) and #11 (need a lot of words to say).
@@
F1: Limitations of Existing Intervention Methods. As noted in the Introduction, current interventions for ADHD fall into two main categories: medical and non-medical. While medical treatments such as Concerta (States, 2023) can have been shown to alleviate communication difficulties (Maras and Hill, 2024), their side effects remain a major concern. As P2 shared, “When I took Concerta, my mind felt clearer, and communication improved. But it also made me tense, nauseous, and disrupted my sleep – so I only take it when absolutely necessary.” Given that adult ADHD is a lifelong condition, the long-term sustainability of medication use is often compromised by these adverse effects. While non-medical interventions were generally perceived as beneficial, participants highlighted their limited applicability in real-world scenarios. Strategies learned in structured environments – such as CBT or SST – often proved difficult to implement during actual conversations due to executive function deficits. As P1 explained, “In real conversations, I just don’t have the mental bandwidth to apply those strategies.” SST, in particular, was critiqued for its lack of ecological validity. C3 noted, “After socializing, individuals with ADHD may overanalyze their mistakes, which increases hesitation and shame in future interactions. Helping them transfer learned skills from training to real-life situations remains a key challenge.”
F2: The Need for Accessible, Therapist-Free, and Real-Time Support. Interviews highlighted a pressing need for accessible, real-time support that operates independently of professional supervision. While ADHD presents persistent communication challenges, its comparatively milder clinical presentation – relative to conditions like depression or mania – often results in less attention and fewer dedicated resources. Only 6 of the 12 participants reported ever receiving structured interventions such as CBT or SST. As P7 noted, “It’s so hard for me to book a therapy appointment. I had to find some books and hope they would help.” This lack of access underscores the value of scalable, automated solutions that provide consistent support without requiring therapist involvement. Participants also emphasized the potential of real-time, context-sensitive assistance. Unlike traditional interventions confined to clinical settings, in-situ support embedded in everyday interactions could bridge the gap between learning and application. As C4 observed, “Providing assistance on-the-fly could turn real-world environments into training spaces, helping ADHD people build resilience and reduce stigma.” Similarly, P3 remarked, “Real-time assistance may help me apply what I’ve been learned immediately.”
@@
In addition to the major themes, several minor findings emerged from the interviews. First, regarding the issue of #12. I have difficulty waiting my turn during conversations, all four clinicians rated this behavior as 3 (frequently observed), and 11 individuals with ADHD reported experiencing similar challenges. However, 10 of them noted that years of social experience had helped them develop self-regulation strategies to manage impulsive behaviors, such as interrupting or speaking out of turn (#12–15). As C4 noted, “This is one of the main differences between adults and children with ADHD. We provide less training on interrupting behaviors for adults.” Additionally, four participants mentioned difficulty maintaining eye contact during conversations, often due to stigma-related self-consciousness. They suggested that supportive tools could help foster more natural communication by encouraging appropriate eye contact.
@@ 3.2.1. Exploratory Brainstorming
To explore a broad range of design ideas for addressing communication challenges related to ADHD, we conducted a group brainstorming session guided by three established principles (Wilson, 2013): (1) focus on quantity, (2) defer judgment, and (3) encourage unconventional thinking. Participants were encouraged to contribute freely without critique, while also building upon each other’s ideas to foster creative synergy. Drawing from the prior interviews, the brainstorming session began with an exploration of commonly reported communication difficulties experienced by individuals with ADHD. Discussions addressed both the limitations of existing interventions and the specific needs for support. Participants were encouraged to generate design concepts for each identified challenge while adhering to the principle of suspending judgment and allowing unrestricted ideation. All ideas proposed during the session were systematically documented. Following the session, researchers and participants collaboratively reviewed and refined the collected concepts, synthesizing a subset of proposals with the greatest potential to address the identified challenges.
@@
[…]
Illustration depicting a storyboard (No. 4) showing the progression of a person’s communication with real-time word suggestions. First panel: A person named Lee looks uneasy saying “um…” while another person appears confused with a question mark, indicating Lee’s struggle to express thoughts, leaving others confused. Second panel: Lee, wearing a virtual reality headset, receives real-time suggestions (represented by ellipsis), helping the conversation get unstuck and flow more smoothly with another person. Third panel: Lee, still with the headset, confidently communicates with flower icons above, indicating his confidence and fluency in conversations with others. Caption for each panel in the image: (1) Lee often gets stuck in conversations, unsure how to express his thoughts. Others are left confused about what he means. (2) With real-time suggestions, Lee gets help just before the conversation gets stuck and continues more smoothly. (3) Lee gradually regains confidence in conversations, and communicates more fluently with others.
@@ 3.2.2. Design Workshop
After formulating initial, high-level design concepts, we refined and expanded these ideas through a dedicated design workshop. Storyboards served as a effective rapid visualization tool (Truong et al., 2006), offering a concrete medium to convey and deepen understanding of abstract design concepts. The workshop featured 18 storyboards (examples shown in Figure 4, with a summary at Figure 11 in Appendix B), each depicting a communication challenge, a proposed solution, or a hypothetical outcome. These storyboards were iteratively developed from the preliminary design ideas, representing concepts at varying levels of granularity.
@@ 3.2.3. Contextual Analysis Findings
F3. MR Devices Support Seamless, Polite Interaction and Real-World Skill Training and Transfer. Participants including P1, P3, and R1 considered laptops and smartphones as impractical for daily communication support, as shifting gaze between the device and conversation partner could be socially inappropriate. While teleprompters were conceptually appealing, their usability was limited by spatial constraints in typical communication settings. Regarding contextual support via ear-worn devices like Memoro (Zulfikar et al., 2024), participants appreciated the reduced visual salience, which may lessen users’ social stigma. However, they expressed concerns about the cognitive demands of processing sequential audio streams. In real-time conversations, managing both the audio streams of partner speech and system prompts was perceived as overwhelming. P2 commented, “As someone with ADHD myself, my brain just … can’t handle that. But combining audio with visual cues might work for me.” Similar concerns about the cognitive load of auditory versus visual cues have been reported in prior work (Villa et al., 2024; Leahy and Sweller, 2016; Klingner et al., 2011).
In contrast, MR devices such as the HoloLens [^3] https://learn.microsoft.com/en-us/hololens/ were favored for their ability to overlay information directly within the user’s field of view, allowing for more natural interactions, such as maintaining eye contact – an essential aspect of effective communication. As P6 remarked, “It’s visually intuitive to present prompts via MR while minimizing conversational disruption. I can engage with my conversation partner while seamlessly receiving information through MR interfaces – without extra effort.” Clinicians also recognized MR’s potential to bridge the gap G1 in sustainable transfer of therapeutic outcomes. As C1 explained, “MR enables SST in real-world scenarios, providing opportunities for ADHD people to learn in context and directly apply their skills in everyday interactions.”
@@ 3.3. Design Goals
[…]
Illustration depicting the interaction flow in Understood, divided into three distinct phases with a clean, structured layout. The Functionality Selection Phase features a horizontal panel with five toggleable icons (Self Summarization, Other Summarization, Word Suggestion, Off-topic Detection) on the left, leading to a prominent “Confirm” button in the center. The Conversation Phase occupies a large central area with a gray field of view, showcasing a trigger element (a small cube) that changes color to indicate on/off-topic status, connected to support panels displaying word suggestions and summaries. The Feedback Phase, on the right, presents a vibrant section with a congratulatory message and colorful confetti icons, all framed within clear, labeled sections and directional arrows guiding the user flow. The caption describes the interaction flow.
@@
[…]
Illustration of the interface for the Auxiliary Functions Selection in Understood, split into two panels labeled A and B. Panel A shows the Functionality Selection Phase with a gray background, featuring five toggleable icons as described in the previous figure. Notably the ”Self-Summarization” button was disabled. Panel B depicts the Conversation Phase interface on a gray background, with a white dashed box indicating the empty space where “Self-Summarization” was removed, alongside a tile for “[Other-Summarization]” listing “Greeting, Repetition, Informal Tone” and another tile for “[Word Suggestions]” offering “Street food, unique flavors.”
@@ 4.1. System Overview
Building on insights from the design workshop, we developed Understood, a MR system designed to support individuals with ADHD in proactively seeking assistance, enhancing communication fluency, and improving conversational performance, which is available on GitHub  [^4] https://github.com/Yu-xinz/UIST2025-8878-Understood. By offering interactive training in realistic settings, the system enables users to gradually refine their speech patterns, serving as a form of automated substitution therapy (DG1). Interaction with Understood is structured into three main phases (Figure 5):
@@ 4.2. Interface and Interaction Design
To accommodate the complexity of MR environments – where real-world backgrounds can introduce distractions, we adopted a minimalist and user-friendly design language. The interface employs a semi-transparent black-and-white color scheme to maintain visual clarity while minimizing cognitive load. In response to potential strain from bright backgrounds, a dark mode was implemented, featuring white text on a black background to reduce visual fatigue. Rounded corners are used consistently throughout the interface to evoke a sense of softness and approachability, ensuring the design remains unobtrusive. To ensure text accessibility, we followed established design guidelines (Williams, 2020; Garnham, 2017; Matsuura et al., 2019) and evaluated several factors, including design consistency, legibility and readability. Based on these criteria, we selected BBC Reith (Maag, 2018) as the system font. Additionally, to mitigate the lack of haptic feedback in MR environments, auditory cues were integrated to provide users with immediate and perceptible confirmation of their actions (DG5). A real-world usage scenario is shown in Figure 7. The individual shown provided informed consent for publication.
Figure 7. An example usage scenario of Understood under a real-world setting.
[…]
Illustration of a real-world usage scenario of Understood, featuring a person sitting at a desk with two computer monitors. The individual is overlaid with translucent interface elements: a “[Self-Summarization]” panel on the left listing “Calendar Update, Package Delivery, John Greeting, Anger Expression,” and an “[Other-Summarization]” panel below noting “Repetitive Greetings, Basic Interaction, Content Deficiency.” A “[Word Suggestions]” panel on the right displays the phrase “I’m not sure.” This demonstrated that Understood’s cues could remain clearly visible against a complicated background.
@@
[…]
@@
[…]
The mapping between each function, its icon and its description.